When 1.0 × 10^12 electrons are transferred from one conductor to another, a potential difference of 10 V appears between the conductors. Calculate the capacitance of the two-conductor system.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. Given that Number of electron = 1 × 10^12 Net charge Q = 1 × 10^12 × 1.6 × 10^–19 = 1.6 × 10^–7 C ∴ The net potential difference = 10 L. ∴ Capacitance – C = q/v = 1.6 * 10^-7/10 = 1.6 * 10^-8 F.